Project

General

Profile

20161129 » History » Version 1

Boyd, Steven, 27 November 2016 14:06

1 1 Boyd, Steven
** MOM report
2
*** Started 'running' on Friday evening with a 24-hour shift.
3
*** ISIS very unstable. Switched on Sunday to TS2 timing.
4
*** Conventional water issues, now resolved.
5
*** Superconducting magnets regularly monitored. Very stable.
6
*** Run plan:
7
**** Have data for 2 of the settings.
8
**** Prepared a third beam setting for recording over night
9
**** More beam tuning tomorrow.
10
*** Loss of shift cover for SS monitoring at 4pm tomorrow.
11
*** Ramp magnets down at 2pm (Agreed).
12
** [[analysis:2016-11-19_physics-shifter|data summary as of 3.30 pm yesterday]] - Rogers gives apologies
13
** SS Magnet status
14
*** Magnet alarms need adjusting, to reflect running performance.
15
*** Feedback from joseph to pierrick regarding operation by non-ss personel.
16
** Hydrogen status
17
*** Cooling, Absorber at 46K. 2nd stage ~26K.
18
*** Monitor overnight and determine coldest temperature. Then will try using heater.
19
** Conventional magnet water flow issue
20
*** Removed PLC from both contentional water pumps.
21
*** Pump set to deliver constrant pressure from configuration app.
22
** Wednesday maintenance plans
23
*** No plans.
24
** Run plans for the next week (discuss)
25
*** Wednesday 23rd : ramp down the SS magnets / maintenance
26
*** Thursday : Shift training + 1 shift - straight track running
27
*** Friday - Sunday : 1 shift per day - straight track running
28
*** Monday  28th : ramp up SS magnets
29
30
* Missing shifts
31
** Still have some missing shift slots for minimal running 
32
*** Friday 25th : BLOC
33
*** Dec 3 : Overnight monitoring shift (00:00 - 08:00)
34
35
* AOB